How many atoms of copper (Cu) are contained in 1.75 moles of copper atoms?

How many moles of aluminum ions (Al 3+) are present in 7.25 x 1025 Al 3+ ions?

How many moles of aspirin contain 8.62 x 10 28 molecules?

How many sugar molecules are contained in 0.475 mole of sugar?

What is the molar mass of nicotine (C10H14N2), the toxic and addictive stimulant in tobacco?

In this balanced chemical equation, 2 C18H30O2 + 49O2 à 36 CO2 + 30 H2O, answer the following questions below:

a. How many atoms of O are found on both sides of the equation? --

b. How many atoms of H are found on both sides of the equation? --

c. What is the mole ratio between O2 and CO2? --

d. What is the mole ratio between C18H30O2 and H2O? --

e. What type of reaction is this?
Answer #1

No of molecules/atoms= no of moles X Avagadro number

1. No of cu atoms= 1.75x6.023x10*23= 10.54 X10*23 atoms

2. No moles= 7.25x10*25/6.023x10*23= 1.2x10*2 moles of Al+3

3. No of moles of asprin= 8.62x10*28/6.023x10*23= 1.43x10*5 moles of asprin

4. No Molecules of sugar= 0.475x6.023x10*23= 2.86 X10*23 sugar molecules